OpenStudy (anonymous):

what is the square root of 200

OpenStudy (alexandervonhumboldt2):

\[\sqrt{200}=\sqrt{100*2}=\sqrt{10*10*2}=10\sqrt{2}\]

OpenStudy (anonymous):

a: 14 b: 20 c: 10/2 d: 5/6

OpenStudy (anonymous):

a

OpenStudy (karolynxx):

C

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what is the square root of 169 - square root of 81

OpenStudy (karolynxx):

13 or 13/2 and 9